A bond has a face value of $1,000 with a maturity date 15 years from today. The bond pays semiannually at a rate of 6% nominal per year based on face value. The interest rate paid on similar corporate bonds has decreased to a current rate of 3%. Determine the market value of the bond.
Answer #1

Coupon payment would be 6/2=3% semi annually or$30 so with I=3% or 1.5% semiannually value of the bond should be,

30/1.015+30/1.015^2+....30/1.015^30+1000/1.015^30=30/1.015(1-(1/1.015^30)/(1-(1/1.015))+1000/1.015^30=$1360.24